# GNN 101: Visual Learning of Graph Neural Networks in Your Web Browser

### v1

@@ 1. Introduction
Graph Neural Networks (GNNs) have become increasingly prominent in deep learning. GNNs offer powerful capabilities for analyzing graph-structured data (e.g., social networks, molecular graphs), which traditional models, mainly designed for non-graph data such as images and text, struggle to handle effectively.
@@
Given the growing popularity of GNNs, numerous educational resources have been developed, covering formats ranging from online blogs to lecture videos and computational notebooks. Despite the great values provided by these educational resources, they have inherent limitations. Specifically, these resources mainly focus on explaining the high-level concepts and model architectures via diagrams or mathematical equations, often failing to present the detailed inner workings of GNNs in an intuitive way.
@@
Interactive visualization can effectively address these limitations, providing an intuitive and engaging approach to learning GNNs. However, visualizing GNNs can be challenging due to the inherent complexity of graph structures and the extensive mathematical details involved in GNN computations. While recent research has proposed visualization systems for GNNs, such as GNNLens (Jin et al., 2022) and CorGIE (Liu et al., 2022), these tools are primarily aimed at AI developers for debugging and improving models. Therefore, they focus on specific aspects related to GNN performances, such as error patterns (Jin et al., 2022) and embedding qualities (Liu et al., 2022), and provide limited support for beginners to learn about GNNs.
@@
In this work, we introduce GNN101, the first interactive visualization tool for learning GNNs. The design of GNN101 is informed by a thorough analysis of existing GNN educational materials. To address the challenges in GNN visualization, GNN101 features a seamless linkage between mathematical formulas and visualizations via different levels of GNN details and provides two complementary views to enhance the learning experience. Specifically, it integrates a model overview, layer operations, and detailed animations for matrix calculations with smooth animations. It complements the advantages of a node-link view, which offers an intuitive representation of the graph structure, and a matrix view, which delivers a space-efficient and comprehensive overview of all features. We are currently deploying GNN101 in three courses that involve the teaching of GNNs at three different universities. This deployment has led to usage scenarios and observational studies that demonstrate the usability of GNN101 and generate design lessons for educational AI visualization tools.
@@
As an open-sourced, web-based tool, GNN101 can be accessed without any installation, ensuring easy and widespread availability. It operates locally in users’ web browsers, supporting a large number of concurrent users. An anonymized version of the source code is available at https://anonymous.4open.science/r/GNN-101, and an anonymous web demo can be accessed at https://gnn101.vercel.app/.
@@ 2. Related Studies
2.1. Visualizations for Algorithms and Machine Learning Models
Using visualization for algorithms can be traced back to the late 1970s (Hundhausen et al., 2002). Previous research has introduced various algorithm visualizations, showcasing both their effectiveness and limitations across different educational contexts (Stasko, 1990; Gurka and Citrin, 1996). However, these studies focused on relatively traditional algorithms and can not be directly applied to visualize the complicated mechanism of modern machine learning models.
With the emerging of machine learning techniques and applications, interactive visualizations of machine learning are gaining wide popularity both in academia and beyond. Visualization techniques have been developed for a variety of deep learning models, including CNNs (Liu et al., 2016; Wang et al., 2020), RNNs (Strobelt et al., 2017; Ming et al., 2017), GANs (Wang et al., 2018), and transformers (Yeh et al., 2023; Bycroft, 2024). However, these techniques are specifically designed for models that work with Euclidean data (e.g., images and text) and cannot be directly applied to GNNs.

### v2

@@ 1. Introduction
Graph Neural Networks (GNNs) offer powerful capabilities for analyzing graph-structured data (e.g., social networks, molecular graphs). This type of data is often difficult to be effectively modeled by traditional machine learning models, which are mainly designed for non-graph data such as images and text. Therefore, GNNs have earned increasing popularity, especially in AI4Science research that leverages the inherent graph structures in scientific data to drive discoveries (e.g., new material design and drug development) (Huang et al., 2024; Wang et al., 2022).
However, learning GNNs can be more challenging than ML models for non-graph data, due to the unique characteristics of graphs and the complexities of graph-based computations. Unlike structured data such as images or tables, graphs are inherently non-Euclidean, meaning their data representation cannot be neatly organized into fixed-size grids. This irregularity makes it difficult to interpret standard computational operations like convolutions or pooling. Meanwhile, graphs can be sparse or have massive sizes, posing significant computational and memory challenges. To address these issues, techniques like graph sampling, mini-batching, and message passing have been developed, which introduce additional complexity for learning GNNs.
@@
Recently, many educational resources of GNNs have been developed, covering formats ranging from online blogs to lecture videos and computational notebooks. Despite the great values provided by these educational resources, they often rely heavily on static illustrations (e.g., diagrams or mathematical equation) for explaining GNNs. Therefore, their focus remains on high-level overviews and theoretical frameworks, with limited emphasis on presenting the detailed inner workings of GNNs in a more intuitive and interactive manner.
@@
Interactive visualization has long been recognized as an effective method for understanding complex algorithms and machine learning models (Wang et al., 2020; Hundhausen et al., 2002). As such, it offers a promising approach to facilitate more effective learning of GNNs by making their intricate processes and concepts more accessible and intuitive. However, visualizing GNNs can be challenging due to the inherent complexity of graph structures and the extensive mathematical details involved in GNN computations. While recent research has proposed visualization systems for GNNs, such as GNNLens (Jin et al., 2022) and CorGIE (Liu et al., 2022), these tools are primarily aimed at AI developers for debugging and improving models. Therefore, they focus on specific aspects related to GNN performances, such as error patterns (Jin et al., 2022) and embedding qualities (Liu et al., 2022), and provide limited support for beginners to learn about GNNs. Most relevant to this paper are the two interactive articles from Distill (Sanchez-Lengeling et al., 2021; Daigavane et al., 2021), which effectively use interactive visualizations to explain various core GNN concepts (e.g., adjacency matrix, convolution, pooling). However, the two papers primarily focus on individual concepts using abstract examples, and do not bridges theoretical understanding with practical application, such as the execution of actual GNN predictions.
@@
To address these limitations and provide a more effective learning experience for GNNs, we introduce GNN101, the first interactive visualization tool for learning GNNs. The design of GNN101 is informed by a thorough analysis of existing GNN educational materials. To address the challenges in GNN visualization, GNN101 features a seamless linkage between mathematical formulas and visualizations via different levels of GNN details and provides two complementary views to enhance the learning experience. Specifically, it integrates a model overview, layer operations, and detailed animations for matrix calculations with smooth animations. It complements the advantages of a node-link view, which offers an intuitive representation of the graph structure, and a matrix view, which delivers a space-efficient and comprehensive overview of all features. We are currently deploying GNN101 in three courses that involve the teaching of GNNs at three different universities. This deployment has led to usage scenarios and observational studies that demonstrate the usability of GNN101 and generate design lessons for educational AI visualization tools.
@@
As an open-sourced, web-based tool, GNN101 can be accessed without any installation, ensuring easy and widespread availability. It operates locally in users’ web browsers, supporting a large number of concurrent users. Source code and an interactive online demo are available at https://github.com/Visual-Intelligence-UMN/GNN-101.
@@ 2. Related Studies
2.1. Visualizations for Learning Algorithms and Machine Learning Models
Interactive visualization has a long history of being used for understanding complicated algorithms and machine learning models. Using visualization for algorithms can be traced back to the 1970s (Hundhausen et al., 2002). By visually representing the dynamic nature of computations (e.g., sorting, searching), visualization facilitates the exploration of algorithmic behavior and enhances the understanding (Stasko, 1990; Gurka and Citrin, 1996). However, these studies focused on more traditional algorithms and may not be directly applied to visualize the complicated mechanism of modern machine learning models.
With the emerging of ML, interactive visualizations of ML are gaining wide popularity both in academia and beyond. Visualization techniques have been developed for a variety of deep learning models, including CNNs (Liu et al., 2016; Wang et al., 2020), RNNs (Strobelt et al., 2017; Ming et al., 2017), GANs (Wang et al., 2018), and transformers (Yeh et al., 2023; Bycroft, 2024). However, these techniques are specifically designed for models that work with Euclidean data (e.g., images and text) and cannot be directly applied to GNNs.
@@ 4.5. Implementation
5. Usage Scenarios
GNN101 is currently being used in three courses that involve the teaching of GNNs at three different universities  [^1] Details omitted for anonymous review. In this section, we show how GNN101 can help a computer science student, Alice, learn various concepts in GNNs based on the reported usage from students.
5.1. Understanding Message Passing
Alice wants to gain a deeper understanding of message passing, the fundamental mechanism underpinning most GNNs (Veličković et al., 2018; Kipf and Welling, 2017; Sanchez-Lengeling et al., 2021). However, the lecture slides and online GNN educational materials primarily explain this concept using diagrams and mathematical formulas, as shown in Figure 1.a, which Alice finds less intuitive and effective. To enhance her learning, Alice turns to GNN101 for assistance.
She begins by exploring the outputs of each layer, which are visualized as a node-link graph, with the features of each node represented as a heatmap. When Alice hovers over a node in the GCNConv2 layer output, GNN101 highlights the selected node, along with its connections to itself and neighboring nodes in the previous layer’s (GCNConv1) output. This interaction helps Alice intuitively grasp the high-level concept of message passing: updating a node’s features by aggregating information from both its neighbors and itself (Figure 2.a).
Curious about the detailed computations of message passing in GCN, Alice clicks on the node to expand the layer and reveal the calculation process as a horizontal flowchart (Figure 2.b). A mathematical formula of the calculation $x^{\prime}_{i}=\sigma(W\sum_{j\in N(i)\cup i}\frac{1}{\sqrt{d_{i}d_{j}}}x_{j}+b)$ also appears to facilitate the understanding. $i$ is the node index, $N(i)$ represents the neighboring nodes of $x_{i}$, $x^{\prime}_{i}$ indicates the updated features for node $i$. On the left of the flowchart, GNN101 shows the relevant node features (heatmaps) from the previous layer, corresponding to $x_{j},j\in N(i)\cup i$ in the formula. The aggregation of these node features $\sum_{j\in N(i)\cup i}\frac{1}{\sqrt{d_{i}d_{j}}}x_{j}$ is visually represented by curves connecting each feature to the aggregation result. The color and thickness of the curves illustrate the weighting factor $\frac{1}{\sqrt{d_{i}d_{j}}}$. The summation result is then multiplied by a weight matrix $W$, with an icon indicating the matrix multiplication, and produces an updated feature. This updated feature is then added with a bias vector $b$ and pass a non-linear activation function $ReLU$, generating the output feature $x_{i}=\sigma(W\sum_{j\in N(i)\cup i}\frac{1}{\sqrt{d_{i}d_{j}}}x_{j}+b)$.
Lastly, to refresh her memory on matrix multiplication and the ReLU activation function, Alice hovers over the results of these calculations, triggering tooltips that explain the process and demonstrate how specific values are computed (Figure 2.D4).
Figure 5. Comparing GNN variants: GNN101 supports the comparisons between GCN (a), GAT (b), and GraphSAGE (c) by visualizing how each model aggregates information from neighboring nodes. Curved edges indicate that parts of the visualization have been omitted from the figure due to space constraints.
5.2. Comparing GNN Variants
After understanding message passing and its implementation in GCN, Alice recalls that message passing is implemented differently across various GNN variants. The lecture slides offer only mathematical formulas for direct comparison of these GNN variants, which limits her ability to fully grasp the differences. Therefore, Alice turns to GNN101to examine other two GNN variants: GAT and GraphSAGE.
Alice starts by selecting GAT in the control panel and clicking on a layer to view its internal computations. The computations are visualized using a horizontal flowchart similar to the one for GCN but with differences in how the input features are connected and aggregated. Unlike GCN, which uses the node degree $\frac{1}{\sqrt{d_{i}d_{j}}}$ as an aggregation factor, GAT computes an attention score $\alpha_{i,j}$, which is visualized with gradient colors for distinction and interactivity. As shown in Figure 5.B2, Alice clicks on an attention score, revealing a step-by-step breakdown of the computation through three mathematical formulas. To aid comprehension, GNN101 includes text annotations and heatmap visualizations to indicate the learnable weight matrix $W$ and the attention $a$ for this layer.
Next, Alice selects GraphSAGE in the control panel and clicks on a layer to view its computations. Sampling icons (Figure 5.C1) appear on the curves connecting and aggregating input features, highlighting the neighbor sampling process, a key innovation of GraphSAGE. Clicking on these icons triggers tooltips that provide textual explanations of the sampling method used in GraphSAGE.
Alice concludes that the primary differences among the three GNN variants lie in how they select and aggregate information from neighbors, which are visually represented by the curves connecting input features to intermediate summation results (Figure 5.a, B1, and c).
Figure 6. Learning different GNN tasks: The final one or two layers handle task-specific processing for graph classification (a), node classification (b), and edge prediction (c). These additional layers also support click-to-expand interactions, similar to the GNN layers. Curved edges indicate that parts of the visualization have been omitted from the figure due to space constraints.
5.3. Learning Different GNN Tasks
After understanding message passing and its implementation across GNN variants, Alice is eager to explore how GNNs solve various tasks in real data. In GNN101, Alice uses the control panel to switch between three common GNN tasks: graph classification, link prediction, and node classification. The matrix view enables Alice to examine the outputs of all layers at the same time (Figure 3.a). She notices that the first several layers of GNNs for these tasks are very similar. Each layer updates the graph node features by learning from its neighbors.
Alice observes the main differences occur in the final one or two layers for task-specific processing (Figure 6). In graph classification (a), a global pooling layer aggregates all node features of the graph and outputs a single feature, which is used to generate the graph classification via an MLP layer. Alice hovers over this layer to see the detailed computation: an element-wise average across dimensions (Figure 4.b). In node classification (b), an MLP layer is applied to each node feature to generate predictions. In edge prediction (c), the features of the two nodes involved in the predicted edge are fed into a prediction layer that outputs a prediction score. Alice clicks on the prediction layer and sees a dot product of the two node features used (C1).
Alice also observes the representation power of GNNs by directly modifying the input data and seeing how a GNN updates the node features layer by layer to generate an accurate prediction.
6. User Interviews
Evaluating educational visualization tools presents challenges due to the long-term nature of learning complex machine learning concepts. Additionally, comparing GNN101 to baselines is difficult because GNN educational visualizations are largely absent. The closest examples are two interactive articles (Sanchez-Lengeling et al., 2021; Daigavane et al., 2021), both of which are lengthy and require over 40 minutes to read even without interacting with their visualizations.  [^2] Based on an average reading speed of 238 words per minute (Brysbaert, 2019). Therefore, we followed practices from previous studies on educational visualizations (Wang et al., 2019, 2020) and conducted an observational study and interviews to assess the usability of GNN101 and to gain insights into user interactions with it.
6.1. Participants
We recruited 12 participants (10 males, 2 females, average age 23.8) through snowball sampling. Given that learning GNNs from scratch is a long-term process, often requiring more time than a typical in-lab interview (30–60 minutes), our study focused on two types of participants: 1) Teaching assistants (TAs) for GNN-related courses who had experience explaining GNN concepts to students in a conversational, interactive format (e.g., office hours, labs, or tutorials). We have 7 participants in this category, denoted as TA1-TA7; 2) Students who had backgrounds in machine learning (average self-reported skill level of 4.5 on a 7-point Likert scale, where 1 represents Novice and 7 represents Expert) but had not formally learned GNNs. We had 5 participants in this category, denoted as S1-S5. None of the participants had prior knowledge of this project.
6.2. Procedure
We conducted the study with participants one-on-one via Zoom. Before the study, each participant completed an intake form that included demographic questions and optional comments on learning GNNs. They also signed a consent form permitting the recording of their audio and screen activity. Each study began with a 5-minute introduction to GNN101 and an overview of its key functions. After the tutorial, participants freely explored GNN101 in their web browsers, following a think-aloud protocol. A manual outlining the main functions of GNN101 was available for reference, and participants could ask questions at any time during the process. Each session concluded with a questionnaire (same as the one used in CNNExplainer (Wang et al., 2020)) and a semi-structured interview. The sessions lasted 30 to 50 minutes, and each participant was compensated with a \$10 Amazon gift card.
6.3. Observations
Figure 7. Results of the post-study questionnaires: (A) User ratings on system usability; (B) User ratings on various components. Ratings were based on a 7-point Likert scale, where 1 indicates strongly disagree and 7 indicates strongly agree.
Overall, participants stated that GNN101 addresses a strong demand for visualization when learning GNNs. For example, TA6 stated “unlike images that look like grids, the unique structure of graph makes it hard to imagine how the data flows through the model… definitely a very useful visualization”. TA5 emphasized the importance of visual aids in GNN learning, stating, “understanding the core message passing steps of GNNs is pretty visual. I used to reference the animation in this distill.pub article (Sanchez-Lengeling et al. (Sanchez-Lengeling et al., 2021)) a lot when explaining things (to students)”. S2 reported “You guys are more interactive, more comprehensive (compared with (Sanchez-Lengeling et al., 2021))…I think it’s pretty helpful, for sure”. TA7 commented “I think this is the best tool out there for learning graph neural network”.
Our observation also showed that GNN101 provides a user-friendly, easy-to-use interface. After the short tutorial, most of the participants (9/12) were able to conduct the free exploration without asking the interviewer questions or referring to the manual. Two participants sought clarification on the visual encoding to confirm their interpretation of the adjacency matrix, which was correct. Even though some participants initially expressed confusion about certain interactions (e.g., saying “Wait, what does this mean?” or repeatedly hovering over specific areas), they were able to resolve these issues independently. For example, after asking a question, TA5 immediately followed by “Oh, I see now, it means the neighbor sampling. That is nice”, before the interviewer had a chance to intervene.
This overall positive attitude is also reflected in the post-study questionnaires, as shown in Figure 7. On a 7-point Likert scale, all average ratings are above or equal to six. Notably, three participants mentioned that they would be continuing as TAs in the upcoming fall semester and asked permission to use GNN101 for the students.
User Interactions: We observed some consistent patterns in how participants interacted with GNN101.
First, most participants (9/12) followed an overview-to-detail approach for exploration: they began by exploring the overall model structure, then selected specific nodes to observe the message passing, and finally hovered over individual components to examine the underlying mathematical operations. Such an interaction flow demonstrates the importance of the hierarchical level of details, as noted by TA4, “I like it … if we present all of the information at once, it might be confusing to know what to focus on first”.
Second, we also observed a tendency to move from intuitive to concrete. All participants chose the node-link view to start their exploration, likely due to its default status and the intuitive representation of graph data. As TA4 commented, “that’s (message passing on graph view) super clear… when I worked as a TA, I found that is where other students have trouble understanding.” Participants then switch to the matrix view to “see the outputs of all layers” and “connect theoretical concept with practical GNN data flows” (TA7).
Third, some participants initially struggled with visualizations but found clarity through the integrated math-visualization linking. For instance, S4 initially felt confused when viewing the message-passing visualization. However, S4 understood the visualization after interacting with the formula, which describes the message passing as the aggregation of neighboring node features. S4 remarked: “This is probably my favorite feature.”
Comparing with Existing Tools: In the post-study interviews, participants frequently compared GNN101 with other GNN educational resources, including interactive articles on distill.pub (Sanchez-Lengeling et al., 2021; Daigavane et al., 2021), lecture slides (Leskovec, 2024), whiteboard drawing, lecture videos, and YouTube educational videos. It’s important to clarify that GNN101is not intended as a complete replacement. For instance, it doesn’t provide the comprehensive textual introductions found in (Sanchez-Lengeling et al., 2021; Daigavane et al., 2021) or the programming implementation details offered by Pytorch Geometric tutorials (Antonio Longa, 2024) (as mentioned by TA5). At the same time, most participants found GNN101 presents a unique advantage, especially through the hierarchical levels of detail and the integration of mathematical formulas with visualization. S3 remarked: “Those are the two things (the hierarchical levels of detail and complementary views) that make it more useful than a YouTube video or the Distill article.”
Suggestions: The most common suggestion we received was to incorporate additional features that support more advanced concepts in GNNs. These suggestions included inductive and transductive learning, the aggregation of both node and edge features, and comparisons among GNNs and other neural network architectures (e.g., MLPs and CNNs). Another frequent suggestion was to allow users to interactively adjust GNN parameters, architecture, or input data and observe the resulting changes. Other suggestions were mainly about minor style changes such as increasing the font size, making the control panel resizable, and adding model training information. These suggestions reflect the participants’ appreciation of the already-implemented functionality and their desire to extend GNN101 to encompass broader concepts. Importantly, even while proposing these enhancements, participants who offered such suggestions agreed that the current version of GNN101 already provided comprehensive support for foundational GNN concepts.